# Universal scaling limits for spin networks via martingale methods

### v1

@@ Abstract
Sapienza Università di Roma, Piazza Aldo Moro 1, 00185 Roma, Italy
In [1, 2, 3] we elaborated a large–deviations theory for disordered systems that allows to study any spin systems by means of simple combinatorial techniques, including the celebrated Sherrington–Kirckpatrick (SK) model, a mean field paradigm for complex systems [4]. Here, we used the same methods to show that any fully connected models with any pairwise interaction, including the random models (i.e., models with random parameters like the SK model), obeys a universal scaling limit. We actually construct the scaling limit of the free energy density in terms of magnetisations eigenstates, i.e., spin states with fixed average magnetisation, in the Eq. (85).
The main points of our analysis where to introduce a sequence of sub–systems of increasing size by partitioning the vertex set into a number of disjoint sub–sets (that in the references [1, 2, 3] we called “layers”) and then by re–joining them up to a certain level of that sequence: this is shown in the Section 1. Then, we reinterpret this as a series of Markov blankets, [5, 6], and identify their Hamiltonian sequence in Section 2. We find that each blanket can be further destructured in two general contributions: the first term, “core” contribution, is the energy from the interactions inside the blanket itself, while the second, the “interface”, is due to the interactions between the spins of the blanket and the rest of the network. Formally, it is the scalar product between the spins of the blanket and local fields that depends on the previous blankets. As we shall see this is a general theory, and up to this point it could be used to deal with any spin system. Then, in Section 3, we show that in fully connected models if the partition is fine enough the core contribution is small, and the thermodynamics is dominated by the interfaces. Finally, in the Section 3.1, we show that the Hamiltonian with the cores removed (IO model, [3]) obeys a universal scaling limit that can be expressed in terms of the magnetisation eigenstates like in the SK model [3, 7].
Apart from providing characterisation for the scaling limit of any fully connected spin network in terms of magnetisation eigenstates, an important goal of this paper is to provide a simple presentation of the method introduced in [1, 2, 3] that is readable by people working in the fields of Bayesian Inference, Machine Learning [5], and Active Inference [6, 8, 9, 10]. Given this purpose, here we avoid the details of the underlying Kernel Theory (KT) [3, 10, 11, 19, 20], within which these results where obtained. The interested reader may look at the Sections 2 and 3 of [3], Section 4 of [10] the SI section of [11] and the preprints [19, 20] for a comprehensive presentation.
@@
and represents one of the possible configurations of the system. Since the support of the single spin is limited to two states, the cardinality of the support of the magnetisations states will be equal to $|\Omega|=2$ raised to the number of vertices $N$,
@@
We also introduce the eigenstates of magnetisation, which we might regard as the fundamental building blocks of our description. We call ‘total magnetisation’ the operator that, if applied to a spin state, returns its total magnetisation. The ‘eigenstates of magnetization’ with eigenvalue $M$ are therefore those spin states with magnetisation exactly equal to $M$. Then, let us introduce the magnetisation function
@@
As for the case of the uniform average, we will use the angle brackets (with the support in subscript) to indicate the average respect to the magnetisation eigenstates,
@@ 1 The general model
this can be rewritten in a way that is convenient for us by introducing the local fields, or cavity fields, depending on the context. By grouping the local fields into a vector
@@
to indicate that the magnetisation state is distributed according to the Gibbs measure $\mu$, which belongs to the set of possible probability measures defined on the support $\Omega^{V}$. The average respect to the generic probability measure will be indicated with angled parenthesis with the measure indicated in subscript, for example:
@@ 1.1 Thermodynamic limit
The very existence of the thermodynamic limit can in fact be reduced to the existence of the limit of $f$ when $N\rightarrow\infty$, which depends critically on how $H_{ij}$ scales in the spin number. The most general form could be site–dependent, but, for simplicity, we will consider a smaller class in which the normalisation is uniform:
$H_{ij}=\frac{\beta J_{ij}}{g\left(V\right)}$ (23)
where $J_{ij}$ are the coupling parameters and $g\left(V\right)$ is a normalisation that ensures the existence of $f$ in the thermodynamic limit, ie. such that the components of $h_{V}$ converges to finite numbers (at least in distribution). For a fully connected model this is usually diverging in the spin number, and not dependent from the specific labels of the spins. For example in the fully connected Curie-Weiss model we find that the correct normalisation must scale with $N$, while for the version with random weights, that is the SK model, the correct normalisation is instead $\sqrt{N}$. Notice that to lighten the notation we reabsorbed also the temperature in the parameters $H_{ij}$.
@@ 1.2 Free energy principle
Applying the Jensen inequality we can easily deduce the canonical free energy principle [6, 8, 10]. Let define the free energy functional
@@
where in the last step we applied the Jensen inequality to the exponential to move the average inside. Then for any $\xi$ holds
@@
that is attained by the Gibbs measure $\mu$. This was the canonical free energy principle. We can also deal with random Hamiltonians. Let now suppose that the coupling matrix is distributed according to some probability measure $\eta$
@@
$N\langle f\rangle_{\eta}=\sup_{\zeta\in\mathscr{P}(\Gamma^{\,V^{2}})}\mathscr{G}\left(\xi\right)$ (33)
@@
that allows to write the free energy in terms of a min-max principle
@@
This is reminiscent of some methods introduced by Guerra et al. and may be related to the min-max principle identified by these authors for the bipartite SK model [15].
@@ 2 Blanket (or Layer) representation
Now that we have introduced the fundamental quantities we can summarise the method proposed in [3]. The theory builds on several Graph–Theoretic methods for statistical physics and probability theory introduced by Lovaz, Borgs, Chayes, Coja-Oghlan, and others, [16, 17, 18, 19, 20]. The idea can be understood also in terms of partitioning the spins into a progression of Markov blankets [5, 6] (that in [1, 2, 3] we called “layers”). For now we will not specify the detailed structure of the blankets, which in general depends on the model, we will just discuss the most general features. Then, let introduce a partition of $V$ into $L$ disjoint subsets, hereafter recalled by the symbol $\mathscr{V}$,
@@
$p:=\left\{p_{\ell}\in\left[0,1\right]:\,\ell\in\Lambda\right\}\in\mathscr{P}(\Lambda).$ (43)
@@
These coefficients are known to Spin Glass theory since they appear in the Parisi functional for the SK model. Notice that $p$ is the discrete canonical conjugate of $q$, and that $p_{\ell}^{2}$ is an infinitesimal quantity that can be ignored in the large $L$ limit. Then, in the large $L$ limit the coefficient converges to the derivative of $q_{\ell}^{2}$ respect to $\ell$.
Figure 1: A comparison between the notion of a Markov blanket as is usually intended (e.g., by Pearl [5] and Friston [6]) versus the ‘layer’ representation of [1, 2, 3]. As one can see from the figures, the second case is just the interaction picture of the first: the blanket is made of vertices while the layer is made of edges. For this paper we will not make the distinction and refer to both as Markov blankets. This is done to facilitate the understanding with other fields that already use this concept.
@@ 2.1 Blankets and conditional expectations
We iterate the recursion from the initial condition $\mathcal{O}_{L}$ until the last step $\mathcal{O}_{0}$ that correspond the thermal average of $\mathcal{O}$ according to $\mu$. Introducing the free energy functional
@@
$Nf=-\log\,\langle\,\exp\,[-Nf\left(\sigma_{V}\right)]\,\rangle_{\mu}.$ (57)
The usual free energy principle is recovered as before by bringing the average respect to $\mu$ inside the exponential, introducing the functional
@@
$N\langle f\rangle_{\eta}=-\langle\,\log\,\langle\,\exp\,[-Nf\left(\sigma_{V}\right)]\rangle_{\mu}\rangle_{\eta},$ (60)
we can bring the average respect to $\eta$ inside the logarithm
@@
$\langle f\rangle_{\eta}\geq\sup_{p\,\in\mathscr{P}(\Lambda)}\mathscr{G}^{*}\left(p\right)$ (62)
We expect that this second variational principle is of the same kind of that appearing in the celebrated Parisi formula for the SK model [4] where, in fact, the prescription is to maximise the the energy functional respect to the order parameter.
@@ 2.2 Blanket internal structure
It has been shown [1] that the blankets possess an internal structure. As one can easily appreciate from Figure 1 of [1], is possible to identify at lest two main types of energy contributions, those corresponding to edges with both ends in $V_{\ell}$, that we call the core, and those with one edge in $V_{\ell}$ and one in $Q_{\ell-1}$, that we call interface:
@@
$\hat{\gamma}_{\ell}:=g\left(V_{\ell}\right)/g\left(V\right),\ \ \ \bar{\gamma}_{\ell}:=g\left(Q_{\ell}\right)/g\left(V\right)$ (64)
Using these coefficients, that for densely connected models can be non–trivial, due to the dependence of $g\left(V\right)$ on the number of spins, the core of the blanket (or layer) is
@@
and is simply a smaller model multiplied by a normalisation. The interface is instead described by a very simple bipartite Hamiltonian, that is the scalar product between the spins of the blanket and the local fields:
@@
Notice that if we interpret $\ell$ as some sort of discrete time, then the the blanket structure is equivalent to the celebrated Transformer circuit [25], with the cores corresponding to the the attention modules [26]. Let conclude by studying the effects of this decomposition on the partition function: to properly explain this it will be convenient to introduce the partition function of the interface,
@@
Then, the average respect to the interface is given by the formula [2]
@@ 3 Fully connected models
In general, for finite dimensional models, or any other networks with irreducible core, we have that the ratio between the core and the interface converges to a constant
@@
and both the cores and the interfaces contribute to the total energy of the blankets in a comparable way also for an infinitely refined partitions, ie. for $L\rightarrow\infty$. In this case the theory is nontrivial and non–mean field. This class includes eg. the Ising model in two and three dimension, polymers [22, 23, 24], random models like the Edaward-Anderson model, but also the Transformer architecture by Google [25, 26] the neural Lattice Field Theory (LFT) in Section 4 of [10], etc.
@@ 3.1 General theory of the IO model
Since the interface seems to have a quite universal scaling limit let construct a general theory of the IO model [3]. We will start from the IO Hamiltonian
@@
If we arrange in decreasing order the field components we can further split $V_{\ell}$ into $L^{\prime}$ sub-blankets with approximately constant value. We call this partition $\mathscr{V}_{\ell}$
@@
Most important, the gap shrinks proportionally to the inverse of $L^{2}$ [3]
@@
$\hat{\gamma}_{\ell\ell^{\prime}}:=g\left(V_{\ell\ell^{\prime}}\right)/g\left(V\right),\ \ \ \hat{h}_{\ell\ell^{\prime}}^{\pm}\left(\sigma_{Q_{\ell-1}}\right):=\hat{\gamma}_{\ell\ell^{\prime}}h_{\ell\ell^{\prime}}^{\pm}\left(\sigma_{Q_{\ell-1}}\right)$ (84)
@@
and can be therefore approximated by energy functions that are supported by the magnetisation eigenstates, similarly to what conjectured by Mezard and Virasoro in [7] for the SK model. This is important, as it allows to work also in the thermodynamic limit using well known methods [28]. Most important, we found that under Gibbs measure the fields are bounded from above/below by the set of parameters [1, 2, 3]
$h^{\pm}\in\mathbb{R}^{\,L^{2}}$ (86)
@@
$\lim_{N\rightarrow\infty}\langle\mathcal{O}\left(\sigma_{V_{\ell\ell^{\prime}}}\right)\rangle_{\bar{\xi}_{\ell\ell^{\prime}}}=\lim_{N\rightarrow\infty}\langle\mathcal{O}(\sigma_{V_{\ell\ell^{\prime}}})\rangle_{m_{\ell\ell^{\prime}}}$ (87)
@@
Finally, notice that due to the new level of partitioning the order parameter evolved from a vector into a matrix, but then, notice that since there is no space structure only the levels matter, and the order parameter is ultimately the distribution of the entries of the parameter matrix.
@@ 3.2 Free energy functional
Now that we have a method to compute the averages it only remains to find a manageable expression for the free energy functional. We start from
@@
$\hat{\gamma}_{i}:=g\left(V_{i}\right)/g\left(V\right),\ \ \ \bar{\gamma}_{i}:=g\left(Q_{i}\right)/g\left(V\right)$ (90)
@@
then if the field is nonrandom we can simply apply the chain rule, while for fields with random coefficients the expression converges to an Ito integral that can be analysed by Ito calculus, see for example [32]. An alternative way to find the free energy functional for the recursion of Eq. (54) is the “Cavity Method” (CM), [4]. The CM is shown in the kernel context in the Section 6 of [3] for the SK model, using a mathematical approach originally developed by Aizenmann, Sims and Starr [33] (see also [1] or [34]).

### v2

@@ Abstract
Sapienza Universit di Roma, Piazza Aldo Moro 1, 00185 Roma, Italy
In [1, 2, 3] we elaborated a Large–Deviations theory for disordered systems that allows to study any spin system by means of standard combinatorial techniques, including the celebrated Sherrington–Kirckpatrick (SK) model, a mean field paradigm for complex systems [4]. Here, we used the same methods to show that any fully connected model with any pairwise interactions, including the random models (i.e., models with random parameters, like the SK model), obeys a universal scaling limit. We actually construct the scaling limit of the free energy density in terms of the magnetisations eigenstates, i.e., spin states with fixed average magnetisation, in the Eq. (85).
The main points of our analysis were: to introduce a sequence of sub–systems of increasing sizes by partitioning the vertex set into a number of disjoint sub–sets (that in the references [1, 2, 3] we called “Layers”) and then by re–joining these up to a certain point of that sequence: this is explained in the Section 1. Then, we reinterpret them as a series of Markov blankets, [5, 6], and identify their Hamiltonian sequence in Section 2. We find that each blanket can be further destructured in two general contributions: the first term, “core” contribution, is the energy from the interactions inside the blanket itself, while the second, the “interface”, is due to the interactions between the spins of the blanket and the rest of the network. Formally, it is the scalar product between the spins of the blanket and local fields that depends on the previous blankets. As we shall see this is a general theory and, up to this point, it could be used to deal with any spin system. Then, in Section 3, we show that in a fully connected model if the partition is refined enough the core contribution is small and the thermodynamics is dominated by the interfaces. Finally, in the Section 3.1, we show that the Hamiltonian with the cores removed (IO model, [3]) obeys a universal scaling limit that can be expressed in terms of the magnetisation eigenstates like in the SK model [3, 7].
Apart from providing characterisation for the scaling limit of any fully connected spin network in terms of magnetisation eigenstates, an important goal of this paper is to provide a simple presentation of the method introduced in [1, 2, 3] that is readable by people working in the fields of Bayesian Inference, Machine Learning [5], and Active Inference [6, 8, 9, 10]. Given this purpose, here we avoid the details of the underlying kernel theory (KT) [3, 10, 11, 19, 20], within which these results where obtained. The interested reader may look at the Sections 2 and 3 of [3], Section 4 of [10], the Theory Insights of [11], and the preprints [19, 20] for a comprehensive presentation.
@@
and represents one of the possible configurations of the system. Since the support of the single spin is limited to two states, the cardinality of the support of the magnetisation states will be equal to $|\Omega|=2$ raised to the number of vertices $N$,
@@
We also introduce the eigenstates of magnetisation, which we might regard as the fundamental building blocks of our description. We call “total magnetisation” the operator that, if applied to a spin state, returns its total magnetisation. The “eigenstates of magnetisation” with eigenvalue $M$ are therefore those spin states with magnetisation exactly equal to $M$. Then, let us introduce the magnetisation function
@@
As for the case of the uniform average, we will use the angle brackets (with the support in subscript) to indicate the average with respect to the magnetisation eigenstates,
@@ 1 The general model
this can be rewritten in a way that is convenient for us by introducing the “local fields”, or “cavity fields”, depending on the context. By grouping the local fields into a vector
@@
to indicate that the magnetisation state is distributed according to the Gibbs measure $\mu$, which belongs to the set of possible probability measures defined on the support $\Omega^{V}$. The average with respect to the generic probability measure will be indicated with angled parenthesis with the measure indicated in subscript, for example:
@@ 1.1 Thermodynamic limit
The very existence of the thermodynamic limit can in fact be reduced to the existence of the limit of $f$ when $N\rightarrow\infty$, which depends critically on how $H_{ij}$ scales in the spin number. The most general form can be site–dependent, but for simplicity we consider a smaller class in which the normalisation is uniform:
$H_{ij}=\frac{\beta J_{ij}}{g\left(V^{2}\right)}$ (23)
here $J_{ij}$ are the coupling parameters and $g(V^{2})$ is a normalisation that ensures the existence of $f$ in the thermodynamic limit, ie. such that the components of $h_{V}$ converges to finite numbers (at least in distribution). For a fully connected model this is usually diverging in the spin number, and not dependent from the specific labels of the spins. For example in the fully connected Curie-Weiss model we find that the correct normalisation must scale with $N$, while for the version with random weights, that is the SK model, the correct normalisation is instead $\sqrt{N}$. Notice that to lighten the notation we reabsorbed also the temperature in the parameters $H_{ij}$.
@@ 1.2 Free energy principle
Applying the Jensen inequality we can easily deduce the canonical free–energy principle [6, 8, 10]. Let define the free energy functional
@@
where in the last step we applied the Jensen inequality to the exponential to bring the average inside. Then for any $\xi$ holds
@@
that is attained by the Gibbs measure, $\mu$. This was the canonical free energy principle. We can also deal with random Hamiltonians. Let now suppose that the coupling matrix is distributed according to some probability measure $\eta$
@@
$N\langle f\rangle_{\eta}=\sup_{\zeta\in\mathscr{P}(\Gamma^{\,V^{2}})}\mathscr{G}\left(\zeta\right)$ (33)
@@
that allows to write the free energy in terms of a min–max principle
@@
This is reminiscent of methods introduced by Guerra et al. and could be related to the min-max principle identified by these authors for the bipartite SK model [15].
@@ 2 Blanket (or Layer) representation
Now that we have introduced the fundamental quantities, we can summarise the method proposed in [1, 2]. The theory builds on several Graph–Theoretical methods for statistical physics and probability theory introduced by Lovaz, Borgs, Chayes, Coja-Oghlan, and others, [16, 17, 18, 19, 20]. The idea can be understood also in terms of partitioning the graph into a progression of Markov blankets [5, 6] (that in [1, 2, 3] where called “layers”). For now we will not specify the detailed structure of the blankets, which in general depends on the model, we just discuss the general features. Then, let introduce a partition of $V$ into $L$ disjoint subsets, hereafter recalled by the symbol $\mathscr{V}$,
@@
$p:=\left\{p_{\ell}\in\left[0,1\right]:\,\ell\in\Lambda\right\}\in\mathscr{P}(\Lambda)$ (43)
@@
These coefficients are known to Spin Glass theory, since they appear in the Parisi functional for the SK model. Notice that $p$ is the discrete canonical conjugate of $q$, and that $p_{\ell}^{2}$ is an infinitesimal quantity that can be ignored in the large $L$ limit. Then, in the large $L$ limit the coefficient converges to the derivative of $q_{\ell}^{2}$ respect to $\ell$.
Figure 1: A comparison between the notion of a Markov blanket as is usually intended (e.g., by Pearl [5] and Friston [6]) versus the “layer” representation of [1, 2, 3]. As one can see from the figures, the second case is just the interaction picture of the first: the blanket is made of vertices while the layer is made of edges. For this paper we will not make the distinction and refer to both as Markov blankets. This is done to facilitate the understanding with other fields that already use this concept.
@@ 2.1 Blankets and conditional expectations
We iterate the recursion from the initial condition $\mathcal{O}_{L}$ until the last step $\mathcal{O}_{0}$ that correspond to the thermal average of $\mathcal{O}$ according to $\mu$. Introducing the free energy functional
@@
$Nf=-\log\,\langle\,\exp\,[-Nf\left(\sigma_{V}\right)]\,\rangle_{\mu},\ \ \ \forall p\in\mathscr{P}(\Lambda).$ (57)
The usual free energy principle is recovered as before by bringing the average with respect to $\mu$ inside the exponential, introducing the functional
@@
$N\langle f\rangle_{\eta}=-\langle\,\log\,\langle\,\exp\,[-Nf\left(\sigma_{V}\right)]\rangle_{\mu}\rangle_{\eta},\ \ \ \forall p\in\mathscr{P}(\Lambda)$ (60)
we can bring the average with respect to $\eta$ inside the logarithm
@@
$N\langle f\rangle_{\eta}\geq\sup_{p\,\in\mathscr{P}(\Lambda)}\mathscr{G}^{*}\left(p\right)$ (62)
We expect that this second variational principle is of the same kind of that appearing in the celebrated Parisi formula for the SK model [4] where, in fact, the prescription is to maximise the energy functional respect to the order parameter.
@@ 2.2 Blanket internal structure
It has been shown in [1] that the blankets posses an internal structure. As one can easily appreciate from Figure 1 of [1], it is possible to identify at least two main types of energy contributions, those corresponding to edges with both ends in $V_{\ell}$, that we call the core, and those with one edge in $V_{\ell}$ and one in $Q_{\ell-1}$, that we call interface:
@@
$\hat{\gamma}_{\ell}:=g\left(V_{\ell}^{2}\right)/g\left(V^{2}\right),\ \ \ \bar{\gamma}_{\ell}:=g\left(\bar{W}_{\ell}\right)/g\left(V^{2}\right)$ (64)
Using these coefficients, that for densely connected models can be non–trivial, due to the dependence of $g(V^{2})$ on the number of spins, the core of the blanket (or layer) is
@@
and is simply a smaller model (multiplied by a normalisation). The interface is instead described by a very simple bipartite Hamiltonian, that is the scalar product between the spins of the blanket and the local fields:
@@
Notice that if we interpret $\ell$ as a sort of discrete time then the blanket structure is reminiscent of the celebrated Transformer architecture [25], with the cores corresponding to the attention modules [26]. Let conclude by studying the effects of this decomposition on the partition function: to properly explain this it will be convenient to introduce the partition function of the interface,
@@
Then, the average with respect to the interface is given by the formula [2]
@@ 3 Fully connected models
In general, for a finite dimensional model, and any other network with irreducible core, we have that the ratio between the core and the interface converges to a constant
@@
hence both the cores and the interfaces contribute to the total energy of the blankets in a comparable way also for an infinitely refined partition, i.e., for $L\rightarrow\infty$. In this case the theory is non–trivial and non–mean field. This class includes, e.g., the Ising model in two and three dimension, polymers [22, 24, 23], random models like the Edaward-Anderson model, but also the Transformer architecture by Google [25, 26], the neural Lattice Field Theory (LFT) in Section 4 of [10], etc.
@@ 3.1 General theory of the IO model
Since the interface seems to posses a quite universal scaling limit, let construct a general theory of the IO model [3]. We will start from the IO Hamiltonian
@@
If we arrange the field components in decreasing order we can further split $V_{\ell}$ into $L^{\prime}$ sub-blankets with approximately constant value. We call this partition $\mathscr{V}_{\ell}$
@@
Most importantly, the gap shrinks proportionally to the inverse of $L^{2}$ [3]
@@
$\hat{\gamma}_{\ell\ell^{\prime}}:=g\left(V_{\ell\ell^{\prime}}^{2}\right)/g\left(V^{2}\right),\ \ \ \hat{h}_{\ell\ell^{\prime}}^{\pm}\left(\sigma_{Q_{\ell-1}}\right):=\hat{\gamma}_{\ell\ell^{\prime}}h_{\ell\ell^{\prime}}^{\pm}\left(\sigma_{Q_{\ell-1}}\right)$ (84)
@@
and can be therefore approximated by energy functions that are supported by the magnetisation eigenstates, similarly to what conjectured Mezard and Virasoro in [7] for the SK model. This is important, as it allows to work also in the thermodynamic limit using well known methods [28]. Most importantly, we found that under Gibbs measure the fields can be approximated by the parameter matrices
$h\in\mathbb{R}^{\,L^{2}},\ \ \ m\in\left[-1,1\right]^{\,L^{2}}$ (86)
@@
$\lim_{N\rightarrow\infty}\langle\mathcal{O}\left(\sigma_{V_{\ell\ell^{\prime}}}\right)\rangle_{\bar{\xi}_{\ell\ell^{\prime}}}=\lim_{N\rightarrow\infty}\langle\mathcal{O}\left(\sigma_{V_{\ell\ell^{\prime}}}\right)\rangle_{m_{\ell\ell^{\prime}}}$ (87)
@@
Finally, notice that due to the new level of partitioning the order parameter evolved from a vector into a matrix, but then, notice that since there is no space structure only the levels matter, and the order parameter is ultimately the distribution of the entries of the parameter matrix, i.e., the quantile $h\left(m\right)$.
@@ 3.2 Free energy functional
Now that we have a method to compute the averages, it only remains to find a manageable expression for the free energy functional. We start from
@@
$\hat{\gamma}_{i}:=g\left(V_{i}^{2}\right)/g\left(V^{2}\right),\ \ \ \bar{\gamma}_{i}:=g\left(W_{i}\right)/g\left(V^{2}\right)$ (90)
@@
then, if the field is non random we can simply apply the chain rule, while for fields with random coefficients the expression converges to an Ito integral, that can be analysed by Ito calculus, see for example [32]. An alternative way to find the free energy functional for the recursion of Eq. (54) is the “Cavity Method” (CM), [4]. The CM is shown in the kernel context in the Section 6 of [3] for the SK model, using a mathematical approach originally developed by Aizenmann, Sims and Starr [33] (see also [1] or [34]).